OpenStudy (anonymous):

in a photoelectric effect experiment, a beryllium surface (work function= 8.01 x 10^-19 J) is bombarded with photons from a 125 nm source of electromagnetic radiation. a.) what is the energy of the incident photons? b.) what is the maximum speed of electrons emitted from the berylliums surface?

OpenStudy (cuanchi):

a) \[E= hc/\lambda \] h= 6.626 x 10^(-34) Js c= 3.0 x 10^8 m/s lambda= 125 x 10^(-9) m

OpenStudy (anonymous):

so maximum speed?

OpenStudy (cuanchi):

with the difference of energy, the energy of the photon - work function you will have the energy of the electron released. you have lambda = h/ mv ; E= hc/lambda v=c/mE v= velocity c= 3.0 x 10^8 m/s m= mass of an electron =9.10938291 x 10^(-31) kilograms E = energy (J) kg m^2/s^2